How many liters of a 15% acid solution should be mixed with 30 liters of a 40% acid solution to obtain a 25% acid solution?
Answer by lostin(9)

Let say we are mixing x litre of solution which is having 15% acid.
So
0.15*x + 0.40* 30 = (x+30)*0.25 ( this equation is mathematical expression of what is written in question. Equation is derived based on the fact that acid quantity will remain same as it was before mixing both the solutions)
0.15*x + 12 = 0.25*x + 7.5
solving this equation
0.10*x = 4.5 ==>>> x = 45 litre